OpenStudy (anonymous):

Important need help ASAP Might give medal/fan How many solutions does the equation have? 1) 2(2x+5)=4(x+3) A. One solution B. Infinite solutions C. No solution 2) a+5=1/5(15+5a) A. One solution B. Infinite solutions C. No solution 3) 4x+1=5+2(2x-4) A. One solution B. Infinite solutions C. No solution Which expressions would complete this equation so that it has one solution? Choose exactly two answers that are correct. 4) 4(2x-3)-4x= A. 4x-12 B. 3x-4 C. 3(2x-6) D. 4(x-5)

OpenStudy (anonymous):

5. What is the best way to classify each equation? 1: 8x+24=2(4x+12) 2: 5x+18-x=2(2x+8) 3: 7(3x-2)=20x-13 4: 3x+2(x-10)=5(x-4) Put A, B, Or C on each number. The same letter can be used several times. A: Identity B: Contradiction C: Neither

OpenStudy (anonymous):

I NEED HELP NOW D:

OpenStudy (anonymous):

Oh no oh no oh no oh no oh no

OpenStudy (anonymous):

@Mathsadness

OpenStudy (anonymous):

@moudown828 PLS ANYBODY THIS IS IMPORTANT

OpenStudy (anonymous):

1 is no solution

OpenStudy (anonymous):

2 is no solution too

OpenStudy (anonymous):

I already had those but thanks alot

OpenStudy (anonymous):

haha sorry. 3 is one solution

OpenStudy (anonymous):

4 is b and c

OpenStudy (anonymous):

and i dont get what 5 is asking

OpenStudy (anonymous):

Its one of those questions where you draw a line from the equations to what classifies them

OpenStudy (anonymous):

i honestly have no clue what that is this is the first time ive ever seen this. sorry. maybe @iambatman can help

OpenStudy (anonymous):

Batman can solve everything :D

OpenStudy (anonymous):

But thanks anyway

OpenStudy (anonymous):

You saved me o_o

OpenStudy (anonymous):

haha im happy to help

OpenStudy (anonymous):

@iambatman I... AM... BATMAN.